OpenStudy (anonymous):

Which of the following points lies on the line whose equation is y = 2x - 3? A.(0,3) B.(1,-1) C.(-1,4)

OpenStudy (smileyxl3):

B.(1,-1)

OpenStudy (anonymous):

substituting (1,-1) i.e. x=1 and y=-1 in the given eq we find LHS= y =-1 RHS= 2x-3= 2*1-3=2-3=-1 Since LHS =RHS therefore B.(1,-1)points lies on the line whose equation is y = 2x - 3.

OpenStudy (anonymous):

@DjNck

OpenStudy (anonymous):

Thanks dpasingh